A 25-ft ladder is placed 8 ft from the base of the house. How high up the side of the house can the ladder reach? Round your answer to the nearest foot.Thanks a lot
Answer by checkley71(8403)

WHAT YOU HAVE IS A RIGHT TRIANGLE WITH AN HYPOTENUSE=25 & 1 SIDE=8.
25^2=8^2+X^2 FOR THE OTHER SIDE OR HIGHT OF THE LADDER.
625=64+X^2
X^2=625-64
X^2=561
X=SQRT561
X=23.68 OR 24 FEET UP THE SIDE OF THE HOUSE.